42.6 ml of KOH was required to titrate 0.137g H2So4. What is the concentration of the solution.
1
Expert's answer
2021-10-18T06:49:09-0400

2KOH + H2SO4 = K2SO4 + 2H2O

According to the equation, 1 mole of H2SO4 requires 2 moles of KOH.

n = m / M

M (H2SO4) = 98.1 g/mol

n (H2SO4) = 0.137 / 98.1 = 0.0014 mol

n (KOH) = 2 x n (H2SO4) = 2 x 0.0014 = 0.0028 mol

CM = n / V

CM (KOH) = 0.0028 / 0.0426 = 0.066 M